A 500g uniform sphere of 7.0cm radius spins at 30 rev/s on an axis through it's centre. Find it's, a) rotational kinetic energy, b)angular momentum, C) radius of gyration
1
Expert's answer
2021-07-16T08:51:38-0400

Given:

"m=0.5\:\rm kg"

"n=30\:\rm rev/s"

"R=0.07\:\rm m"

(a)

"E_k=\frac{I\omega^2}{2}=\frac{2/5mR^2*(2\pi n)^2}{2}=4/5\pi^2mR^2n^2"

"=4/5\pi^2*0.5*0.07^2*30^2=17.4\:\rm J"

(b)

"L=I\omega=2/5mR^2*2\pi n=4/5\pi mR^2n"

"=4/5\pi*0.5*0.07^2*30=0.37\:\rm kg\cdot m^2/s"

(c)

"r=\sqrt{\frac{I}{m}}=\sqrt{\frac{2R^2}{5}}"

"=\sqrt{2/5}*0.07=0.044\:\rm m=4.4\:\rm cm"